### v2

# Present address: Physics Dept., U.C. Santa Barbara, Santa Barbara, California 93106, USA and Google Quantum AI, 301 Mentor Dr, Goleta, CA 93111, USA

Abstract
Ionizing radiation impacts create bursts of quasiparticle density in superconducting qubits. These bursts temporarily degrade qubit coherence which can be detrimental for quantum error correction. Here, we experimentally resolve quasiparticle bursts in 3D gap-engineered transmon qubits by continuously monitoring qubit transitions. Gap engineering allowed us to reduce the burst detection rate by a factor of a few. This reduction falls several orders of magnitude short of that expected if the quasiparticles were to quickly thermalize to the cryostat temperature. We associate the limited effect of gap engineering with the slow thermalization of the phonons in our chips after the burst.
preprint: APS/123-QED
Introduction.—Quantum error correction protocols redundantly encode a logical qubit in the entangled state of multiple physical qubits [nielsen_quantum_2010], making encoded information robust to errors on individual qubits. This holds if errors are uncorrelated between different physical qubits and lack temporal correlations on the same qubit. Common errors in superconducting processors (e.g. qubit relaxation from material defects) generally satisfy these assumptions. However, there are rare error burst events that violate them. During these error bursts (hereby referred to as “bursts”), qubit relaxation rates temporarily increase across multiple qubits simultaneously [vepsalainen_impact_2020, martinis_saving_2021, wilen_correlated_2021, mcewen_resolving_2022, thorbeck_two-level-system_2023, harrington_synchronous_2024, li_direct_2024]. Mitigating bursts is therefore essential to realize large-scale quantum error correction.
The underlying mechanism of the bursts is the impact of ionizing radiation on the chip hosting the qubits  [^1] There is another mechanism for bursts, where a phonon shower is generated by the avalanche-like release of stress in the chip [yelton_correlated_2025]. Its experimental signatures and mitigation strategies should be similar to those for high-energy impacts.. A high-energy particle striking the chip produces a cascade of high-energy phonons that rapidly propagate throughout the substrate. When the phonons collide with superconducting films of the qubits, they break Cooper pairs, producing quasiparticle excitations (QPs). When the QPs tunnel across the Josephson junction, they can exchange energy with the qubit degree of freedom. This results in an elevated qubit relaxation rate. Only when the QPs recombine or get trapped does the error burst end and the processor performance recover. This prompts the search for strategies to either accelerate the phonon decay [martinis_saving_2021, iaia_phonon_2022] or suppress QP tunneling.
On the latter front, a promising way forward is engineering of the superconducting gap [aumentado_nonequilibrium_2004, ferguson_microsecond_2006, naaman_narrow-band_2007, court_energy_2007, shaw_kinetics_2008, sun_measurements_2012, riwar_efficient_2019, serniak_nonequilibrium_2019, diamond_distinguishing_2022, pan_engineering_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024, mcewen_resisting_2024]. A non-uniform gap profile, which can be created by varying superconducting film thicknesses, can impede QP tunneling across the qubit junction and suppress decoherence [diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024, mcewen_resisting_2024]. For this approach to be effective, the difference in gaps at the two sides of the junction should exceed the qubit energy by an amount significantly larger than the QP energy. This condition has been shown to hold for the “resident” non-equilibrium QPs typically present in the superconducting films in the steady state [connolly_coexistence_2024]. By interacting with phonons, the energy of these QPs equilibrates with the base plate of the refrigerator even though their density remains strongly out of equilibrium. Yet, an important question for the suppression of QP-induced correlated errors remains unanswered: can gap engineering prevent QP tunneling following the impact of high-energy particles, when the phonon bath itself is moved away from equilibrium?
In this paper, we address this question by measuring QP bursts in 3D transmons with different pairs of junction film thicknesses and, consequently, with different gap profiles. First, we benchmark our gap engineering by measuring the gap difference, $\delta\Delta$, for our devices with different film thicknesses. The gap difference is obtained by monitoring the tunneling rate of resident QPs at different fridge temperatures. By varying the film thickness, we control the gap difference from 0 to 10 $\mathrm{GHz}$ (in frequency units). We then evaluate the effect of gap difference on the detection rate of burst events and measure qubit relaxation rates during bursts.
We find that gap engineering reduces the burst detection rate by a factor of a few. However, the reduction falls short of the four orders of magnitude suppression observed for the resident QPs. We believe that this discrepancy results from the elevation of the effective chip temperature during the burst to about 90 mK. The additional thermal energy allows the QPs to overcome the gap difference and tunnel across the junction, thus facilitating qubit relaxation. We directly confirm the persistence of elevated substrate temperature by comparing the qubit excitation and relaxation rates during the burst.
Figure 1: (a) Parity-switching rate ($\Gamma$) as a function of the fridge temperature ($T$) for devices with different film thicknesses. The film thickness is used as a knob to control the gap difference $\>\delta\Delta$ at the qubit junction. Down (up)-pointing triangles represent rate $\Gamma_{0}$ ($\Gamma_{1}$) conditioned on the qubit residing in the ground (excited) state. When $\>\delta\Delta>hf_{q}$, the gap difference is efficient in obstructing the tunneling of resident QPs. Correspondingly, $\Gamma_{1}$ follows an Arrhenius law with the activation exponent $\>\delta\Delta-hf_{q}$, i.e., $\Gamma_{1}\propto\exp\left(-\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T\right)$. This regime is realized in the big- and medium-$\>\delta\Delta$ device (left and middle panels). For the small-$\>\delta\Delta$ device (right panel), $\>\delta\Delta<hf_{q}$, and the gap difference is not effective at preventing the tunneling of resident QPs. The observed rate $\Gamma_{1}$ for the small-$\>\delta\Delta$ device is the highest among the measured devices. Comparing the data for $\Gamma_{1}$ and $\Gamma_{0}$ to theory (solid lines) allows us to extract $\>\delta\Delta$ and the resident QP density for all three devices. (b) Exponential suppression of the resident QP tunneling rate, $\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}}$, in state $|1\rangle$. Different points correspond to devices with different $\>\delta\Delta-\>hf_{q}$, different cooldowns, and temperatures [noauthor_see_nodate]. For convenience of comparison between different devices, the rate is normalized by the resident QP density, $\>\gamma^{\mathrm{QP}}_{1}\coloneqq(\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}})/\>x_{\mathrm{QP}}^{\mathrm{ne}}$. When the gap difference surpasses the qubit energy, the normalized tunneling rate decreases exponentially with increasing $\>(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$. The inset shows the full dependence of $\>\gamma^{\mathrm{QP}}_{1}$ on $\>(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$. The normalized rate diverges when $\>\delta\Delta=\>hf_{q}$ but becomes temperature-insensitive for $\>\delta\Delta<\>hf_{q}$. The gray-shaded regions in the main plot and inset represent the theoretical prediction for the rate for a range of additional parameters such as $E_{J}$ or $E_{C}$.
Our results should be contrasted with Ref. [mcewen_resisting_2024], where the gap difference – obtained by film thicknesses similar to our gap-engineered device – resulted in essentially no bursts being detected. The difference may arise from a combination of several factors. First, unlike ours, the chip in Ref. [mcewen_resisting_2024] has most of its area covered by the aluminum ground plane. Second, their substrate is silicon whereas ours is sapphire. Third, the chips are thermalized differently. Our chips are anchored to the package only by two isolated clamps each, whereas Ref. [mcewen_resisting_2024] presumably used numerous wirebonds that anchor their chip. Finally, unlike in our measurement, Ref. [mcewen_resisting_2024] probed the qubit only sparsely, once every 100 $\mathrm{\mu s}$. Such a measurement may miss bursts recovering faster than 100 $\mathrm{\mu s}$.
Experimental setup.— Our goal is to evaluate the effectiveness of gap engineering in suppressing decoherence caused by the QP bursts. The gap difference between the junction leads, with gaps of $\Delta$ and $\Delta+\>\delta\Delta$, prevents the tunneling of low-energy QPs from the low-gap lead to the high-gap one. The elevation of the superconducting gap in one of the leads, $\delta\Delta$, is controlled by changing the film thicknesses employed in the junction fabrication [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024]. We fabricate three 3D aluminum transmons using different film thickness pairs: [17 nm, 83 nm], [23 nm, 123 nm], and [85 nm, 106 nm] (measured by atomic force microscopy). Based on prior measurements [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, marchegiani_quasiparticles_2022], the corresponding $\>\delta\Delta$ are expected to be approximately 10 GHz (“big”-$\>\delta\Delta$), 5 GHz (“medium”-$\>\delta\Delta$), and 0.5 GHz (“small”-$\>\delta\Delta$), respectively, in frequency units. Table 1 summarizes relevant qubit parameters.
The transmons are designed to be in the offset-charge-sensitive regime [riste_millisecond_2013, serniak_hot_2018, serniak_direct_2019], with the charge dispersion on the order of MHz. This enables the detection of charge-parity-switching events, which are experimental signatures of QP tunneling. By measuring the parity-switching rates, we can monitor how frequently QPs cross the junction in the presence of $\>\delta\Delta$. All three chips are mounted together in a multiplexed package, ensuring that electromagnetic environment is the same [noauthor_see_nodate].
Table 1: Fit parameters and uncertainties corresponding to devices in Fig. 1(a). For the medium- and small-$\>\delta\Delta$ devices, the parity-switching rate is dominated by the QP tunneling rate, making it difficult to extract $\Gamma_{1}^{\mathrm{ph}}$. Therefore, for these devices, we assume $\Gamma_{0}^{\mathrm{ph}}=\Gamma_{1}^{\mathrm{ph}}$. In contrast, we can confidently fit $\Gamma_{1}^{\mathrm{ph}}$ for the big-$\>\delta\Delta$ device. This value can be found in [noauthor_see_nodate].
$f_{q}$ $\>\delta\Delta/h$ $\Delta/h$ $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ $\Gamma_{0}^{\mathrm{ph}}$
(GHz) (GHz) (GHz) (×$\times 10^{-10}$) ($\mathrm{s}^{-1}$)
big-$\>\delta\Delta$ 3.57 8.2$\pm$0.4 45.2$\pm$0.4 6.3$\pm$1.8 0.22$\pm$0.01
med.-$\>\delta\Delta$ 4.78 5.8$\pm$0.1 46.1$\pm$0.1 1.0$\pm$0.1 0.24$\pm$0.01
small-$\>\delta\Delta$ 4.27 0$\pm$1.1 45.1$\pm$0.4 1.9$\pm$0.1 0.52$\pm$0.05
Measuring the gap difference.—As an initial step, we characterize the superconducting gap barriers, $\delta\Delta$, of our devices. To do this, we measure the parity-switching rates in the steady state, i.e., away from rare QP bursts. In this case, the parity-switching rates are dominated by the tunneling of resident non-equilibrium QPs populating superconducting films. The energy distribution of these QPs equilibrates with the base stage of the refrigerator [connolly_coexistence_2024]. The temperature dependence of the QP tunneling rate then follows the Arrhenius law, with the activation exponent controlled by $\>\delta\Delta$. Indeed, $\>\delta\Delta$ impedes the QPs across the junction, such that only few energetic QPs tunnel through the junction. Fitting the measured temperature dependence of the parity-switching rates to the Arrhenius law thus directly yields $\>\delta\Delta$, as we explain below. This measurement resembles that in Ref. [connolly_coexistence_2024]. Figure 1(a) shows the measured parity-switching rates as a function of temperature for our three devices.
To fit the measured rates, we use a detailed theoretical model of QP tunneling in the presence of a gap difference at the junction [diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024]. According to this theory, when the qubit is in the ground state, the parity-switching rate reads $\>\Gamma^{\mathrm{QP}}_{0}=\>\gamma^{\mathrm{QP}}_{0}\>x_{\mathrm{QP}}^{\mathrm{ne}}$. Here, $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ is the temperature-independent density of resident QPs (in units of Cooper pair density), and $\>\gamma^{\mathrm{QP}}_{0}\propto E_{J}\exp\left(-\>\delta\Delta/\>k_{\mathrm{B}}T\right)$ is proportional to the tunneling rate and the fraction of QPs with energy exceeding $\delta\Delta$. Here, $\>k_{\mathrm{B}}$ is the Boltzmann constant, $T$ is the QP temperature, and $E_{J}$ is the Josephson energy of the junction. When the qubit is excited, and if $\>\delta\Delta>\>hf_{q}$, this expression modifies to $\>\Gamma^{\mathrm{QP}}_{1}=\>\gamma^{\mathrm{QP}}_{1}\>x_{\mathrm{QP}}^{\mathrm{ne}}$ with $\>\gamma^{\mathrm{QP}}_{1}\propto E_{J}\exp\left(-\left[\>\delta\Delta-\>hf_{q}\right]/\>k_{\mathrm{B}}T\right)$, where $h$ is the Planck constant and $f_{q}$ is the qubit frequency. The activation energy is reduced compared to that in $\>\gamma^{\mathrm{QP}}_{0}$ because the QP can absorb the excitation from the qubit, making it easier to overcome the gap difference. On contrary, if $\>\delta\Delta<\>hf_{q}$, then $\>\gamma^{\mathrm{QP}}_{1}$ becomes roughly temperature-independent. For each device, we simultaneously fit $\Gamma_{0}$ and $\Gamma_{1}$ to the expressions outlined above, using $\>\delta\Delta$ and $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ as fitting parameters.
There are two complications in this fitting protocol. First, at high temperatures, $T>100\>\mathrm{mK}$, the density of equilibrium QPs, $\>x_{\mathrm{QP}}^{\mathrm{th}}(T)=\sqrt{2\pi\>k_{\mathrm{B}}T/\Delta}\exp\left(-\Delta/\>k_{\mathrm{B}}T\right)$, becomes comparable to that of the resident non-equilibrium QPs, $\>x_{\mathrm{QP}}^{\mathrm{ne}}$. Therefore, at high temperatures, we account for the contribution of equilibrium QPs to the parity-switching rates. Second, there is a small background contribution $\Gamma^{\mathrm{ph}}$ to parity-switching rates coming from the absorption of stray pair-breaking photons at the junction [houzet_photon-assisted_2019, diamond_distinguishing_2022, connolly_coexistence_2024]. In our theoretical fit, we add such contributions as a temperature-independent offset. Notably, the contribution $\Gamma^{\mathrm{ph}}$ of the photons might differ for the two qubit states [houzet_photon-assisted_2019, diamond_distinguishing_2022]. We use notations $\Gamma_{0}^{\mathrm{ph}}$ and $\Gamma_{1}^{\mathrm{ph}}$ to differentiate this contribution between the respective qubit states.
Figure 2: (a) Schematic of how a high-energy impact leads to a QP burst. The collision of high-energy particles ionizes electron-hole pairs in the substrate. When the electron-hole pairs recombine, they generate high-energy phonons (green waves). These phonons have enough energy to break Cooper pairs in superconducting films, thereby generating numerous QPs. Tunneling of these QPs across the junction increases the qubit relaxation rate. (b) Pulse sequence for detecting QP bursts. Every 5.7 $\mathrm{\mu s}$, the qubit is read-out, then projected into |$\ket{1}$. (c) Histogram of the qubit relaxation event counts within a 1 ms window. Different colors correspond to different devices. In the absence of bursts, we expect the histograms to represent a Poisson process. The most likely number of relaxation events is determined by the steady-state coherence time $T_{1}^{\rm steady}$. The positions of the observed maxima are consistent with measured steady-state relaxation times ($T_{1}^{\rm steady}\approx 100\mathrm{\mu s}$). We associate clear deviation from the Poissonian statistics in the tails of the distributions with the rare QP bursts. To the right of the black line, the measured counts for both devices are dominated by QP bursts. Post-selecting events based on this threshold thus allows fair comparison of the burst occurrence rate in the big-$\>\delta\Delta$ and small-$\>\delta\Delta$ devices. The larger $\>\delta\Delta$ yields roughly five times fewer detected burst events according to this criteria. (d) A readout trace during a QP burst. After the burst begins, the qubit experiences an excess number of relaxation events. During the burst, the number of relaxation events in several 1 ms bins exceeds the threshold in panel (c). It allows us to identify the burst. (e) Excess qubit relaxation rate $\Delta\Gamma_{10}$ after high-energy impacts, averaged over all observed burst events. The bursts are identified according to the threshold defined in panel (c). Exponential fit results are shown with solid lines. The extracted burst duration, see the legend, is insensitive to $\>\delta\Delta$. The inset shows the qubit temperature, $T_{q}$, during bursts in the big-$\>\delta\Delta$ device (see [noauthor_see_nodate] for the measurement protocol). We observe $T_{q}$ abruptly increases from the background value (50 mK) to $\approx 90$ mK and stays consistently elevated for over 5 ms. We attribute this increase to elevated QP temperature during the burst. We note that the 50 mK background likely does not reflect the real temperature of the sample. Excess of |$\ket{1}$ measurements primarily stems from measurement errors.
Burst measurement.—With the gap differences under control, we proceed to analyze the temporally correlated errors caused by QP bursts in the devices. To assess the effect of the gap difference on mitigating the high-energy impacts, we compare the bursts in the big- and small-$\>\delta\Delta$ devices. The results for the medium-$\>\delta\Delta$ device are discussed separately in [noauthor_see_nodate], due to the complications arising from its gap difference being comparable to the qubit energy.
To detect the QP bursts, we attempt to stabilize the qubit in its excited state. To this end, we repeatedly readout the qubit state every 5.7 $\mathrm{\mu s}$. If we find the qubit in its ground state $|0\rangle$, we apply a $\pi$-pulse to bring it to the excited state $|1\rangle$. The pulse sequence is illustrated in Fig. 2(b). During QP burst events, the qubit relaxation rate significantly increases. This allows us to identify the bursts by the excess number of observed $|1\rangle\rightarrow|0\rangle$ relaxation events [e.g. Fig. 2(d)]. We note that the measurement is post-selected for events not causing large changes to offset charge [noauthor_see_nodate]. However, we find that this post-selection is inconsequential for the burst event statistics.
To collect sufficient statistics, we repeat the measurement for about 10 hours for each device. We then chop the measurement traces into $1\>\mathrm{ms}$ windows and tally the number of observed relaxation events in each window. Normalized histograms of the relaxation counts are shown in Fig. 2(c). For each device, the histogram contains a large peak (10-15 relaxation events). This peak is unrelated to the QP bursts. It describes a Poissonian relaxation process consistent with the independently measured steady-state qubit relaxation time $T_{1}^{\rm steady}\sim 100\>\mu\mathrm{s}$. In addition to the Poissonian peak, however, there is an anomalous tail with an excess number of relaxation events, which stems from the bursts. Its clear distinction from the Poissonian peak makes us confident that a burst has occurred within a $1\>\mathrm{ms}$ window when the number of relaxation events exceeded a certain threshold  [^2] To increase the confidence in identifying the bursts, we additionally rely on temporal correlation of the relaxation events and reject sequences where the qubit has leaked to more highly excited states, see [noauthor_see_nodate] for details. Choosing a common threshold for both devices, we see that the burst detection rate in the big-$\delta\Delta$ device ($0.57\>\mathrm{min}^{-1}$) is lower than in the small-$\>\delta\Delta$ device ($2.76\>\mathrm{min}^{-1}$).
Notably, the suppression of the burst detection rate by gap engineering is only about a factor of five. This observation is inconsistent with the QP energy thermalizing with the mixing chamber of the fridge immediately after the impact. If QPs were thermalized – as for resident QPs in Fig. 1 – $\>\delta\Delta$ would suppress the burst detection rate by about a factor of $\exp(\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T)\sim 10^{4}$ in the big-$\>\delta\Delta$ device. Here, $T=25\>\mathrm{mK}$ is the base temperature of our fridge.
Next, we consider how $T_{1}$ of the qubit returns to its steady state after the impact. To this end, we use continuous monitoring data from the measurement of Fig. 2(b-d). We identify the bursts using the threshold shown in Fig. 2(c). Averaging the measurement outcomes between many observed burst events allows us to determine the qubit relaxation rate $\Gamma_{10}$ as a function of time after the impact [noauthor_see_nodate]. Fig. 2(e) shows the results. For both devices, the relaxation rate returns to the steady state rougly exponentially. The associated time constants $\sim 0.7\>\mathrm{ms}$ are insensitive to the gap engineering.
In addition to the qubit relaxation rate $\Gamma_{10}(t)$ after the impact, we similarly measure the qubit excitation rate $\Gamma_{01}(t)$. To this end, we repeat the measurement similar to that described in Fig. 2(b), except without resetting the qubit to |$\ket{1}$. From the two rates, one may extract the time dependence of the QP density, $\>x_{\mathrm{QP}}(t)$, and the qubit temperature $T_{q}(t)$, see the inset of Fig. 2(e) for the latter quantity. Here, the qubit temperature is defined as $\>k_{\mathrm{B}}T_{q}=hf_{q}/\mathrm{ln}\left(\Gamma_{10}/\Gamma_{01}\right)$. The time dependence of $\>x_{\mathrm{QP}}$ is similar to that of $\Gamma_{10}(t)$. In about a millisecond, the QP density decays to a level where QPs no longer limit the qubit $T_{1}$.
The qubit temperature $T_{q}$ surges to about 90 mK after the impact, significantly exceeding the fridge’s base temperature. The qubit temperature returns to its steady state exponentially, with a time constant of about $6\>\mathrm{ms}$. This time scale is almost 10 times longer than the $0.7\>\mathrm{ms}$ time constant for the decay of $\>x_{\mathrm{QP}}$. The temperature thus remains elevated even long after the generated by the impact non-equilibrium QPs are gone. This points to a dominant role of the substrate properties in determining the evolution of $T_{q}(t)$. Indeed, the observed increase in qubit temperature to $90\>\mathrm{mK}$ is consistent with the substrate phonon heat capacity for a typical radiation impact energy of $\sim 100\mathrm{keV}-1\>\mathrm{MeV}$ [fowler_spectroscopic_2024, noauthor_see_nodate]. The decay constant for the qubit temperature is determined by the rate of phonon escape from the system, likely through the chip clamps.
The elevated substrate temperature after burst explains the weak sensitivity of burst occurrence to the gap difference. As previously described, the suppression of QP-induced qubit relaxation by $\>\delta\Delta$ is controlled by the activation exponent $\exp(-[\delta\Delta-hf_{q}]/\>k_{\mathrm{B}}T)$. For the gap difference to be effective, $\delta\Delta-hf_{q}$ must be much larger than $\>k_{\mathrm{B}}T$. For our big-$\delta\Delta$ device, $(\delta\Delta-hf_{q})/\>k_{\mathrm{B}}\sim 200\>\mathrm{mK}$ which only exceeds the lattice temperature during the QP burst by a factor of two. This is consistent with the observed decrease of the burst occurrence by about a factor of a few in the presence of gap difference.
Conclusion.— In this work, we investigate the effect of superconducting gap engineering on QP bursts in transmon qubits. Our results confirm that the superconducting gap difference between the two sides of the junction exceeding the qubit energy, $\>\delta\Delta>hf_{q}$, can reduce both the burst detection rate, see Fig. 2(c), and the qubit relaxation rate during these events, see Fig. 2(e).
However, we also uncover a critical limitation of this method: during bursts, the substrate and QPs retain high energy $\sim 90\>\mathrm{mK}$, only mildly lower than $\>\delta\Delta-hf_{q}$. We attribute this to the slow escape of hot phonons from our chip. Our findings suggest two potential strategies to further suppress the adverse effects of bursts. First, increasing the gap difference to $\>\delta\Delta/\>k_{\mathrm{B}}\gg 100\>\mathrm{mK}$ can prevent QPs from crossing the junction even when their energy is elevated. The increase of $\>\delta\Delta$ can be achieved with high-gap materials, such as granular aluminum [abeles_enhancement_1966, cohen_superconductivity_1968, deutscher_transition_1973, sun_measurements_2012]. Second, the effectiveness of gap engineering can be enhanced by providing an efficient escape mechanism for hot phonons. This involves improving the thermal link between the chip and the fridge’s base plate. Alternatively, the phonons can be absorbed by patches of normal metal placed directly on the chip [martinis_saving_2021, iaia_phonon_2022, larson_quasiparticle_2025]. Due to their high heat capacity, these patches can serve as efficient heat sinks. Combining these strategies should suppress the detrimental effect of QP bursts on superconducting quantum processors.
Acknowledgements.— We acknowledge Suhas Ganjam who provided package and qubit design. We acknowledge useful discussions with Valla Fatemi and Kaavya Sahay. This research was sponsored by the Army Research Office (ARO) under grants no.W911NF-22-1- 0053 and W911NF-23-1-0051, by DARPA under grant no. HR0011-24-2-0346, by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DARPA, DoE, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. Fabrication facilities use was supported by the Yale Institute for Nanoscience and Quantum Engineering (YINQE) and the Yale Univeristy Cleanroom.

### v3

@@
 Here, we experimentally resolve quasiparticle bursts in 3D gap-engineered transmon qubits by continuously monitoring qubit transitions.
-Gap engineering allowed us to reduce the burst detection rate by a factor of a few.
-This reduction falls several orders of magnitude short of that expected if the quasiparticles were to quickly thermalize to the cryostat temperature.
+Gap engineering allows us to reduce the burst detection rate by a factor of five.
+This reduction falls four orders of magnitude short of that expected if the quasiparticles were to quickly thermalize to the cryostat temperature.
 We associate the limited effect of gap engineering with the slow thermalization of the phonons in our chips after the burst.
@@
 However, there are rare error burst events that violate them.
-During these error bursts (hereby referred to as “bursts”), qubit relaxation rates temporarily increase across multiple qubits simultaneously [vepsalainen_impact_2020, martinis_saving_2021, wilen_correlated_2021, mcewen_resolving_2022, thorbeck_two-level-system_2023, harrington_synchronous_2024, li_direct_2024].
+During these error bursts (hereby referred to as “bursts”), qubit relaxation rates temporarily increase across multiple qubits simultaneously [vepsalainen_impact_2020, martinis_saving_2021, wilen_correlated_2021, mcewen_resolving_2022, thorbeck_two-level-system_2023, harrington_synchronous_2025, li_cosmic-ray-induced_2025].
 Mitigating bursts is therefore essential to realize large-scale quantum error correction.
@@
 Its experimental signatures and mitigation strategies should be similar to those for high-energy impacts..
-A high-energy particle striking the chip produces a cascade of high-energy phonons that rapidly propagate throughout the substrate.
+As illustrated in Fig.
+2(a), a high-energy particle striking the chip produces a cascade of high-energy phonons that rapidly propagate throughout the substrate.
 When the phonons collide with superconducting films of the qubits, they break Cooper pairs, producing quasiparticle excitations (QPs).
@@
 Yet, an important question for the suppression of QP-induced correlated errors remains unanswered: can gap engineering prevent QP tunneling following the impact of high-energy particles, when the phonon bath itself is moved away from equilibrium?
-In this paper, we address this question by measuring QP bursts in 3D transmons with different pairs of junction film thicknesses and, consequently, with different gap profiles.
+In this paper, we address this question by measuring QP bursts in transmons with different pairs of junction film thicknesses and, consequently, with different gap profiles.
+All of our transmons are made of aluminum, have 3D geometry [paik_observation_2011], and are deposited on sapphire substrates.
 First, we benchmark our gap engineering by measuring the gap difference, $\delta\Delta$, for our devices with different film thicknesses.
@@
 We then evaluate the effect of gap difference on the detection rate of burst events and measure qubit relaxation rates during bursts.
-We find that gap engineering reduces the burst detection rate by a factor of a few.
+We find that gap engineering reduces the burst detection rate by a factor of five.
 However, the reduction falls short of the four orders of magnitude suppression observed for the resident QPs.
@@
 Figure 1: (a) Parity-switching rate ($\Gamma$) as a function of the fridge temperature ($T$) for devices with different film thicknesses.
-The film thickness is used as a knob to control the gap difference $\>\delta\Delta$ at the qubit junction.
+The film thickness is used as a knob to control the gap difference $\delta\Delta$ at the qubit junction.
 Down (up)-pointing triangles represent rate $\Gamma_{0}$ ($\Gamma_{1}$) conditioned on the qubit residing in the ground (excited) state.
-When $\>\delta\Delta>hf_{q}$, the gap difference is efficient in obstructing the tunneling of resident QPs.
-Correspondingly, $\Gamma_{1}$ follows an Arrhenius law with the activation exponent $\>\delta\Delta-hf_{q}$, i.e., $\Gamma_{1}\propto\exp\left(-\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T\right)$.
-This regime is realized in the big- and medium-$\>\delta\Delta$ device (left and middle panels).
-For the small-$\>\delta\Delta$ device (right panel), $\>\delta\Delta<hf_{q}$, and the gap difference is not effective at preventing the tunneling of resident QPs.
-The observed rate $\Gamma_{1}$ for the small-$\>\delta\Delta$ device is the highest among the measured devices.
-Comparing the data for $\Gamma_{1}$ and $\Gamma_{0}$ to theory (solid lines) allows us to extract $\>\delta\Delta$ and the resident QP density for all three devices.
+When $\delta\Delta>hf_{q}$, the gap difference is efficient in obstructing the tunneling of resident QPs.
+Correspondingly, $\Gamma_{1}$ follows an Arrhenius law with the activation exponent $\delta\Delta-hf_{q}$, i.e., $\Gamma_{1}\propto\exp\left(-\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T\right)$.
+This regime is realized in the big- and medium-$\delta\Delta$ device (left and middle panels).
+For the small-$\delta\Delta$ device (right panel), $\delta\Delta<hf_{q}$, and the gap difference is not effective at preventing the tunneling of resident QPs.
+The observed rate $\Gamma_{1}$ for the small-$\delta\Delta$ device is the highest among the measured devices.
+Comparing the data for $\Gamma_{1}$ and $\Gamma_{0}$ to theory (solid lines) allows us to extract $\delta\Delta$ and the resident QP density for all three devices.
+The resulting fit parameters are shown in Table 1.
+Notably, $\Gamma^{\mathrm{ph}}$ and $x_{\mathrm{QP}}^{\mathrm{ne}}$ for the three devices are of the same order of magnitude despite being fabricated individually.
+We attribute this consistency to the shared electromagnetic and radiation environment provided by the multiplexed package [noauthor_see_nodate].
 (b) Exponential suppression of the resident QP tunneling rate, $\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}}$, in state $|1\rangle$.
-Different points correspond to devices with different $\>\delta\Delta-\>hf_{q}$, different cooldowns, and temperatures [noauthor_see_nodate].
-For convenience of comparison between different devices, the rate is normalized by the resident QP density, $\>\gamma^{\mathrm{QP}}_{1}\coloneqq(\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}})/\>x_{\mathrm{QP}}^{\mathrm{ne}}$.
-When the gap difference surpasses the qubit energy, the normalized tunneling rate decreases exponentially with increasing $\>(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$.
-The inset shows the full dependence of $\>\gamma^{\mathrm{QP}}_{1}$ on $\>(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$.
-The normalized rate diverges when $\>\delta\Delta=\>hf_{q}$ but becomes temperature-insensitive for $\>\delta\Delta<\>hf_{q}$.
+Different points correspond to devices with different $\delta\Delta-\>hf_{q}$, different cooldowns, and temperatures [noauthor_see_nodate].
+For convenience of comparison between different devices, the rate is normalized by the resident QP density, $\gamma^{\mathrm{QP}}_{1}\coloneqq(\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}})/\>x_{\mathrm{QP}}^{\mathrm{ne}}$.
+When the gap difference surpasses the qubit energy, the normalized tunneling rate decreases exponentially with increasing $(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$.
+The inset shows the full dependence of $\gamma^{\mathrm{QP}}_{1}$ on $(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$.
+The normalized rate diverges when $\delta\Delta=\>hf_{q}$ but becomes temperature-insensitive for $\delta\Delta<\>hf_{q}$.
 The gray-shaded regions in the main plot and inset represent the theoretical prediction for the rate for a range of additional parameters such as $E_{J}$ or $E_{C}$.
@@
 Third, the chips are thermalized differently.
-Our chips are anchored to the package only by two isolated clamps each, whereas Ref.
+Our chips are anchored to the package only by two isolated clamps each [noauthor_see_nodate], whereas Ref.
 [mcewen_resisting_2024] presumably used numerous wirebonds that anchor their chip.
@@
 We fabricate three 3D aluminum transmons using different film thickness pairs: [17 nm, 83 nm], [23 nm, 123 nm], and [85 nm, 106 nm] (measured by atomic force microscopy).
-Based on prior measurements [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, marchegiani_quasiparticles_2022], the corresponding $\>\delta\Delta$ are expected to be approximately 10 GHz (“big”-$\>\delta\Delta$), 5 GHz (“medium”-$\>\delta\Delta$), and 0.5 GHz (“small”-$\>\delta\Delta$), respectively, in frequency units.
+Based on prior measurements [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, marchegiani_quasiparticles_2022], the corresponding $\delta\Delta$ are expected to be approximately 10 GHz (“big”-$\delta\Delta$), 5 GHz (“medium”-$\delta\Delta$), and 0.5 GHz (“small”-$\delta\Delta$), respectively, in frequency units.
 Table 1 summarizes relevant qubit parameters.
@@
 This enables the detection of charge-parity-switching events, which are experimental signatures of QP tunneling.
-By measuring the parity-switching rates, we can monitor how frequently QPs cross the junction in the presence of $\>\delta\Delta$.
+By measuring the parity-switching rates, we can monitor how frequently QPs cross the junction in the presence of $\delta\Delta$.
 All three chips are mounted together in a multiplexed package, ensuring that electromagnetic environment is the same [noauthor_see_nodate].
@@
 1(a).
-For the medium- and small-$\>\delta\Delta$ devices, the parity-switching rate is dominated by the QP tunneling rate, making it difficult to extract $\Gamma_{1}^{\mathrm{ph}}$.
+For the medium- and small-$\delta\Delta$ devices, the parity-switching rate is dominated by the QP tunneling rate, making it difficult to extract $\Gamma_{1}^{\mathrm{ph}}$.
 Therefore, for these devices, we assume $\Gamma_{0}^{\mathrm{ph}}=\Gamma_{1}^{\mathrm{ph}}$.
-In contrast, we can confidently fit $\Gamma_{1}^{\mathrm{ph}}$ for the big-$\>\delta\Delta$ device.
+In contrast, we can confidently fit $\Gamma_{1}^{\mathrm{ph}}$ for the big-$\delta\Delta$ device.
 This value can be found in [noauthor_see_nodate].
-$f_{q}$ $\>\delta\Delta/h$ $\Delta/h$ $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ $\Gamma_{0}^{\mathrm{ph}}$
-(GHz) (GHz) (GHz) (×$\times 10^{-10}$) ($\mathrm{s}^{-1}$)
-big-$\>\delta\Delta$ 3.57 8.2$\pm$0.4 45.2$\pm$0.4 6.3$\pm$1.8 0.22$\pm$0.01
-med.-$\>\delta\Delta$ 4.78 5.8$\pm$0.1 46.1$\pm$0.1 1.0$\pm$0.1 0.24$\pm$0.01
-small-$\>\delta\Delta$ 4.27 0$\pm$1.1 45.1$\pm$0.4 1.9$\pm$0.1 0.52$\pm$0.05
+$f_{q}$ $\delta\Delta/h$ $\Delta/h$ $x_{\mathrm{QP}}^{\mathrm{ne}}$ $\Gamma_{0}^{\mathrm{ph}}$
+(GHz) (GHz) (GHz) ($\times 10^{-10}$) ($\mathrm{s}^{-1}$)
+[…]
+[…]
+[…]
 Measuring the gap difference.—As an initial step, we characterize the superconducting gap barriers, $\delta\Delta$, of our devices.
@@
 The energy distribution of these QPs equilibrates with the base stage of the refrigerator [connolly_coexistence_2024].
-The temperature dependence of the QP tunneling rate then follows the Arrhenius law, with the activation exponent controlled by $\>\delta\Delta$.
-Indeed, $\>\delta\Delta$ impedes the QPs across the junction, such that only few energetic QPs tunnel through the junction.
-Fitting the measured temperature dependence of the parity-switching rates to the Arrhenius law thus directly yields $\>\delta\Delta$, as we explain below.
+The temperature dependence of the QP tunneling rate then follows the Arrhenius law, with the activation exponent controlled by $\delta\Delta$.
+Indeed, $\delta\Delta$ impedes the QPs across the junction, such that only few energetic QPs tunnel through the junction.
+Fitting the measured temperature dependence of the parity-switching rates to the Arrhenius law thus directly yields $\delta\Delta$, as we explain below.
 This measurement resembles that in Ref.
 [connolly_coexistence_2024].
-Figure 1(a) shows the measured parity-switching rates as a function of temperature for our three devices.
+Fig.
+1(a) shows the measured parity-switching rates as a function of temperature for our three devices, and Fig.
+1(b) shows a summary of additional devices and cooldowns.
 To fit the measured rates, we use a detailed theoretical model of QP tunneling in the presence of a gap difference at the junction [diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024].
-According to this theory, when the qubit is in the ground state, the parity-switching rate reads $\>\Gamma^{\mathrm{QP}}_{0}=\>\gamma^{\mathrm{QP}}_{0}\>x_{\mathrm{QP}}^{\mathrm{ne}}$.
-Here, $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ is the temperature-independent density of resident QPs (in units of Cooper pair density), and $\>\gamma^{\mathrm{QP}}_{0}\propto E_{J}\exp\left(-\>\delta\Delta/\>k_{\mathrm{B}}T\right)$ is proportional to the tunneling rate and the fraction of QPs with energy exceeding $\delta\Delta$.
-Here, $\>k_{\mathrm{B}}$ is the Boltzmann constant, $T$ is the QP temperature, and $E_{J}$ is the Josephson energy of the junction.
-When the qubit is excited, and if $\>\delta\Delta>\>hf_{q}$, this expression modifies to $\>\Gamma^{\mathrm{QP}}_{1}=\>\gamma^{\mathrm{QP}}_{1}\>x_{\mathrm{QP}}^{\mathrm{ne}}$ with $\>\gamma^{\mathrm{QP}}_{1}\propto E_{J}\exp\left(-\left[\>\delta\Delta-\>hf_{q}\right]/\>k_{\mathrm{B}}T\right)$, where $h$ is the Planck constant and $f_{q}$ is the qubit frequency.
-The activation energy is reduced compared to that in $\>\gamma^{\mathrm{QP}}_{0}$ because the QP can absorb the excitation from the qubit, making it easier to overcome the gap difference.
-On contrary, if $\>\delta\Delta<\>hf_{q}$, then $\>\gamma^{\mathrm{QP}}_{1}$ becomes roughly temperature-independent.
-For each device, we simultaneously fit $\Gamma_{0}$ and $\Gamma_{1}$ to the expressions outlined above, using $\>\delta\Delta$ and $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ as fitting parameters.
+According to this theory, when the qubit is in the ground state, the parity-switching rate reads $\Gamma^{\mathrm{QP}}_{0}=\>\gamma^{\mathrm{QP}}_{0}\>x_{\mathrm{QP}}^{\mathrm{ne}}$.
+Here, $x_{\mathrm{QP}}^{\mathrm{ne}}$ is the temperature-independent density of resident QPs (in units of Cooper pair density), and $\gamma^{\mathrm{QP}}_{0}\propto E_{J}\exp\left(-\>\delta\Delta/\>k_{\mathrm{B}}T\right)$ is proportional to the tunneling rate and the fraction of QPs with energy exceeding $\delta\Delta$.
+Here, $k_{\mathrm{B}}$ is the Boltzmann constant, $T$ is the QP temperature, and $E_{J}$ is the Josephson energy of the junction.
+When the qubit is excited, and if $\delta\Delta>\>hf_{q}$, this expression modifies to $\Gamma^{\mathrm{QP}}_{1}=\>\gamma^{\mathrm{QP}}_{1}\>x_{\mathrm{QP}}^{\mathrm{ne}}$ with $\gamma^{\mathrm{QP}}_{1}\propto E_{J}\exp\left(-\left[\>\delta\Delta-\>hf_{q}\right]/\>k_{\mathrm{B}}T\right)$, where $h$ is the Planck constant and $f_{q}$ is the qubit frequency.
+The activation energy is reduced compared to that in $\gamma^{\mathrm{QP}}_{0}$ because the QP can absorb the excitation from the qubit, making it easier to overcome the gap difference.
+On contrary, if $\delta\Delta<\>hf_{q}$, then $\gamma^{\mathrm{QP}}_{1}$ becomes roughly temperature-independent.
+For each device, we simultaneously fit $\Gamma_{0}$ and $\Gamma_{1}$ to the expressions outlined above, using $\delta\Delta$ and $x_{\mathrm{QP}}^{\mathrm{ne}}$ as fitting parameters.
 There are two complications in this fitting protocol.
-First, at high temperatures, $T>100\>\mathrm{mK}$, the density of equilibrium QPs, $\>x_{\mathrm{QP}}^{\mathrm{th}}(T)=\sqrt{2\pi\>k_{\mathrm{B}}T/\Delta}\exp\left(-\Delta/\>k_{\mathrm{B}}T\right)$, becomes comparable to that of the resident non-equilibrium QPs, $\>x_{\mathrm{QP}}^{\mathrm{ne}}$.
+First, at high temperatures, $T>100\>\mathrm{mK}$, the density of equilibrium QPs, $x_{\mathrm{QP}}^{\mathrm{th}}(T)=\sqrt{2\pi\>k_{\mathrm{B}}T/\Delta}\exp\left(-\Delta/\>k_{\mathrm{B}}T\right)$, becomes comparable to that of the resident non-equilibrium QPs, $x_{\mathrm{QP}}^{\mathrm{ne}}$.
 Therefore, at high temperatures, we account for the contribution of equilibrium QPs to the parity-switching rates.
@@
 We use notations $\Gamma_{0}^{\mathrm{ph}}$ and $\Gamma_{1}^{\mathrm{ph}}$ to differentiate this contribution between the respective qubit states.
-Figure 2: (a) Schematic of how a high-energy impact leads to a QP burst.
-The collision of high-energy particles ionizes electron-hole pairs in the substrate.
-When the electron-hole pairs recombine, they generate high-energy phonons (green waves).
-These phonons have enough energy to break Cooper pairs in superconducting films, thereby generating numerous QPs.
-Tunneling of these QPs across the junction increases the qubit relaxation rate.
-(b) Pulse sequence for detecting QP bursts.
-Every 5.7 $\mathrm{\mu s}$, the qubit is read-out, then projected into |$\ket{1}$.
-(c) Histogram of the qubit relaxation event counts within a 1 ms window.
-Different colors correspond to different devices.
-In the absence of bursts, we expect the histograms to represent a Poisson process.
-The most likely number of relaxation events is determined by the steady-state coherence time $T_{1}^{\rm steady}$.
-The positions of the observed maxima are consistent with measured steady-state relaxation times ($T_{1}^{\rm steady}\approx 100\mathrm{\mu s}$).
-We associate clear deviation from the Poissonian statistics in the tails of the distributions with the rare QP bursts.
-To the right of the black line, the measured counts for both devices are dominated by QP bursts.
-Post-selecting events based on this threshold thus allows fair comparison of the burst occurrence rate in the big-$\>\delta\Delta$ and small-$\>\delta\Delta$ devices.
-The larger $\>\delta\Delta$ yields roughly five times fewer detected burst events according to this criteria.
-(d) A readout trace during a QP burst.
-After the burst begins, the qubit experiences an excess number of relaxation events.
-During the burst, the number of relaxation events in several 1 ms bins exceeds the threshold in panel (c).
-It allows us to identify the burst.
-(e) Excess qubit relaxation rate $\Delta\Gamma_{10}$ after high-energy impacts, averaged over all observed burst events.
-The bursts are identified according to the threshold defined in panel (c).
-Exponential fit results are shown with solid lines.
-The extracted burst duration, see the legend, is insensitive to $\>\delta\Delta$.
-The inset shows the qubit temperature, $T_{q}$, during bursts in the big-$\>\delta\Delta$ device (see [noauthor_see_nodate] for the measurement protocol).
-We observe $T_{q}$ abruptly increases from the background value (50 mK) to $\approx 90$ mK and stays consistently elevated for over 5 ms.
-We attribute this increase to elevated QP temperature during the burst.
-We note that the 50 mK background likely does not reflect the real temperature of the sample.
-Excess of |$\ket{1}$ measurements primarily stems from measurement errors.
+[…]
 Burst measurement.—With the gap differences under control, we proceed to analyze the temporally correlated errors caused by QP bursts in the devices.
-To assess the effect of the gap difference on mitigating the high-energy impacts, we compare the bursts in the big- and small-$\>\delta\Delta$ devices.
-The results for the medium-$\>\delta\Delta$ device are discussed separately in [noauthor_see_nodate], due to the complications arising from its gap difference being comparable to the qubit energy.
-To detect the QP bursts, we attempt to stabilize the qubit in its excited state.
+To assess the effect of the gap difference on mitigating the high-energy impacts, we compare the bursts in the big- and small-$\delta\Delta$ devices.
+The results for the medium-$\delta\Delta$ device are discussed separately in [noauthor_see_nodate], due to the complications arising from its gap difference being comparable to the qubit energy.
+To detect the QP bursts, we attempt to stabilize the qubit in its excited state [mcewen_resolving_2022, harrington_synchronous_2025, dominicis_evaluating_2024, wu_mitigating_2025].
 To this end, we repeatedly readout the qubit state every 5.7 $\mathrm{\mu s}$.
@@
 2(d)].
-We note that the measurement is post-selected for events not causing large changes to offset charge [noauthor_see_nodate].
-However, we find that this post-selection is inconsequential for the burst event statistics.
+In our measurement, we post-select events that do not cause large changes to the offset charge.
+However, we find that this post-selection is inconsequential for the burst event statistics [noauthor_see_nodate].
 To collect sufficient statistics, we repeat the measurement for about 10 hours for each device.
@@
 2(c).
-For each device, the histogram contains a large peak (10-15 relaxation events).
-This peak is unrelated to the QP bursts.
-It describes a Poissonian relaxation process consistent with the independently measured steady-state qubit relaxation time $T_{1}^{\rm steady}\sim 100\>\mu\mathrm{s}$.
-In addition to the Poissonian peak, however, there is an anomalous tail with an excess number of relaxation events, which stems from the bursts.
-Its clear distinction from the Poissonian peak makes us confident that a burst has occurred within a $1\>\mathrm{ms}$ window when the number of relaxation events exceeded a certain threshold  [^2] To increase the confidence in identifying the bursts, we additionally rely on temporal correlation of the relaxation events and reject sequences where the qubit has leaked to more highly excited states, see [noauthor_see_nodate] for details.
-Choosing a common threshold for both devices, we see that the burst detection rate in the big-$\delta\Delta$ device ($0.57\>\mathrm{min}^{-1}$) is lower than in the small-$\>\delta\Delta$ device ($2.76\>\mathrm{min}^{-1}$).
+For each device, the histogram contains a large peak (10 - 15 relaxation events), which is unrelated to the QP bursts.
+Although the peak of the distribution is described by the uncorrelated Poissonian relaxation events, there is an anomolous non-Poissonian tail containing an excess number of relaxation events.
+This tail stems from bursts containing correlated errors.
+Its clear distinction from the Poissonian peak makes us confident that a burst has occurred within a $1\>\mathrm{ms}$ window when the number of relaxation events exceeded a certain threshold.
+In practice, we set the threshold at eight standard deviations away from the mean of both Poissonian distributions to make the false positive event rate negligible.
+Notably, some of the identified events are unrelated to bursts; they stem from unstable lossy defects in the environment [muller_interacting_2015, klimov_fluctuations_2018, burnett_decoherence_2019, muller_towards_2019, schlor_correlating_2019, lisenfeld_electric_2019].
+These events can be effectively filtered out since their time dynamics is sufficiently different from that of the bursts [noauthor_see_nodate].
+Using the common threshold for both devices, we see that the burst detection rate in the big-$\delta\Delta$ device ($0.57\>\mathrm{min}^{-1}$) is lower than that in the small-$\delta\Delta$ device ($2.76\>\mathrm{min}^{-1}$).
 Notably, the suppression of the burst detection rate by gap engineering is only about a factor of five.
@@
 If QPs were thermalized – as for resident QPs in Fig.
-1 – $\>\delta\Delta$ would suppress the burst detection rate by about a factor of $\exp(\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T)\sim 10^{4}$ in the big-$\>\delta\Delta$ device.
+1 – $\delta\Delta$ would suppress the burst detection rate by about a factor of $\exp(\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T)\sim 10^{4}$ in the big-$\delta\Delta$ device.
 Here, $T=25\>\mathrm{mK}$ is the base temperature of our fridge.
@@
 Averaging the measurement outcomes between many observed burst events allows us to determine the qubit relaxation rate $\Gamma_{10}$ as a function of time after the impact [noauthor_see_nodate].
-Fig.
-2(e) shows the results.
-For both devices, the relaxation rate returns to the steady state rougly exponentially.
+The results of this analysis are shown in Fig.
+2(e).
+For both devices, the return of the relaxation rate to the steady state is rougly exponential.
 The associated time constants $\sim 0.7\>\mathrm{ms}$ are insensitive to the gap engineering.
-In addition to the qubit relaxation rate $\Gamma_{10}(t)$ after the impact, we similarly measure the qubit excitation rate $\Gamma_{01}(t)$.
-To this end, we repeat the measurement similar to that described in Fig.
-2(b), except without resetting the qubit to |$\ket{1}$.
-From the two rates, one may extract the time dependence of the QP density, $\>x_{\mathrm{QP}}(t)$, and the qubit temperature $T_{q}(t)$, see the inset of Fig.
-2(e) for the latter quantity.
-Here, the qubit temperature is defined as $\>k_{\mathrm{B}}T_{q}=hf_{q}/\mathrm{ln}\left(\Gamma_{10}/\Gamma_{01}\right)$.
-The time dependence of $\>x_{\mathrm{QP}}$ is similar to that of $\Gamma_{10}(t)$.
-In about a millisecond, the QP density decays to a level where QPs no longer limit the qubit $T_{1}$.
+We note that different ionizing radiation sources, such as muons or gamma rays, may produce bursts of varying severity since vastly different amounts of energy can be absorbed by the chip [fowler_spectroscopic_2024].
+Our measurements lack resolution to distinguish between different ionizing radiation sources.
+This is the reason why we average the response of all bursts together, regardless of their severity or origin.
+In the supplementary materials, we provide evidence that up to an overall rescaling, the dynamics of the qubit’s relaxation rate is independent of the burst severity [noauthor_see_nodate].
+We leave a detailed investigation of different radiation sources to future work [vepsalainen_impact_2020, harrington_synchronous_2025, li_cosmic-ray-induced_2025, fowler_spectroscopic_2024].
+[…]
 The qubit temperature $T_{q}$ surges to about 90 mK after the impact, significantly exceeding the fridge’s base temperature.
 The qubit temperature returns to its steady state exponentially, with a time constant of about $6\>\mathrm{ms}$.
-This time scale is almost 10 times longer than the $0.7\>\mathrm{ms}$ time constant for the decay of $\>x_{\mathrm{QP}}$.
+This time scale is almost 10 times longer than the $0.7\>\mathrm{ms}$ time constant for the decay of $x_{\mathrm{QP}}$.
 The temperature thus remains elevated even long after the generated by the impact non-equilibrium QPs are gone.
@@
 The elevated substrate temperature after burst explains the weak sensitivity of burst occurrence to the gap difference.
-As previously described, the suppression of QP-induced qubit relaxation by $\>\delta\Delta$ is controlled by the activation exponent $\exp(-[\delta\Delta-hf_{q}]/\>k_{\mathrm{B}}T)$.
-For the gap difference to be effective, $\delta\Delta-hf_{q}$ must be much larger than $\>k_{\mathrm{B}}T$.
+As previously described, the suppression of QP-induced qubit relaxation by $\delta\Delta$ is controlled by the activation exponent $\exp(-[\delta\Delta-hf_{q}]/\>k_{\mathrm{B}}T)$.
+For the gap difference to be effective, $\delta\Delta-hf_{q}$ must be much larger than $k_{\mathrm{B}}T$.
 For our big-$\delta\Delta$ device, $(\delta\Delta-hf_{q})/\>k_{\mathrm{B}}\sim 200\>\mathrm{mK}$ which only exceeds the lattice temperature during the QP burst by a factor of two.
-This is consistent with the observed decrease of the burst occurrence by about a factor of a few in the presence of gap difference.
+This is consistent with the observed five-fold decrease of the burst occurrence in the presence of gap difference.
+At short times ($t\lesssim 100\>\mu\mathrm{s}$), $\Delta\Gamma_{10}(t)$ in the big-$\delta\Delta$ device deviates from exponential decrease, as shown in Fig.
+2(e).
+We attribute this to the time dynamics of the QP distribution temperature.
+For $t\lesssim 100\>\mu\mathrm{s}$, the effective qubit temperature $T_{q}$ spikes well above 100 mK before rapidly decaying to $\sim 90$ mK for the remainder of the burst ($100\>\mu\mathrm{s}<t<5\>\mathrm{ms}$) [Fig.
+2(e) inset and Fig.
+S14(f) in supplementary materials].
+This temperature spike is likely related to the abundance of high-energy phonons that quickly get down-converted [kurilovich_correlated_2025].
+Combined with the strong temperature dependence of $\gamma^{\mathrm{QP}}_{10}$ in big-$\delta\Delta$ device, the rapid spike and decay of the temperature is responsible for the non-exponential behavior of $\Delta\Gamma_{10}(t)$ during the initial part of the burst.
 Conclusion.— In this work, we investigate the effect of superconducting gap engineering on QP bursts in transmon qubits.
-Our results confirm that the superconducting gap difference between the two sides of the junction exceeding the qubit energy, $\>\delta\Delta>hf_{q}$, can reduce both the burst detection rate, see Fig.
+Our results confirm that the superconducting gap difference between the two sides of the junction exceeding the qubit energy, $\delta\Delta>hf_{q}$, can reduce both the burst detection rate, see Fig.
 2(c), and the qubit relaxation rate during these events, see Fig.
 2(e).
-However, we also uncover a critical limitation of this method: during bursts, the substrate and QPs retain high energy $\sim 90\>\mathrm{mK}$, only mildly lower than $\>\delta\Delta-hf_{q}$.
+However, we also uncover a critical limitation of this method: during bursts, the substrate and QPs retain high energy $\sim 90\>\mathrm{mK}$, only mildly lower than $\delta\Delta-hf_{q}$.
 We attribute this to the slow escape of hot phonons from our chip.
 Our findings suggest two potential strategies to further suppress the adverse effects of bursts.
-First, increasing the gap difference to $\>\delta\Delta/\>k_{\mathrm{B}}\gg 100\>\mathrm{mK}$ can prevent QPs from crossing the junction even when their energy is elevated.
-The increase of $\>\delta\Delta$ can be achieved with high-gap materials, such as granular aluminum [abeles_enhancement_1966, cohen_superconductivity_1968, deutscher_transition_1973, sun_measurements_2012].
+First, increasing the gap difference to $\delta\Delta/\>k_{\mathrm{B}}\gg 100\>\mathrm{mK}$ can prevent QPs from crossing the junction even when their energy is elevated.
+The increase of $\delta\Delta$ can be achieved with high-gap materials, such as granular aluminum [abeles_enhancement_1966, cohen_superconductivity_1968, deutscher_transition_1973, sun_measurements_2012].
 Second, the effectiveness of gap engineering can be enhanced by providing an efficient escape mechanism for hot phonons.
 This involves improving the thermal link between the chip and the fridge’s base plate.
+To achieve optimal thermalization, one should test various substrate materials and clamping schemes; this comparison is outside the scope of our work.
 Alternatively, the phonons can be absorbed by patches of normal metal placed directly on the chip [martinis_saving_2021, iaia_phonon_2022, larson_quasiparticle_2025].